OpenStudy (anonymous):

Solve the linear system using the method of substituition x + y = -2 y = x + 6

hero (hero):

x + y = -2 y = x + 6 y = -2 - x y = x + 6 y = y -2 - x = x + 6 2x = -8 x = -4 y = 2 The intersection point is (-4,2).

OpenStudy (anonymous):

Thanks